(-3,-3) m=1/2 Find the equation of the line through the given point with the given slope.
This is what I did and is this correct??
y-(-3)=1/2(x-(-3))
y+3=1/2x+6
-3 -3
y=1/2x+3

Answer by rfer(16322)

y=mx+b
-3=0.5*-3+b
-3=-1.5+b
-3+1.5=b
-1.5=b
y=0.5x-1.5